Use the Pythagorean Theorem to find the length from L to M to the nearest tenth.
L has the coordinates (2, -8) and M has the coordinates (9, 15)

Answer by Boreal(15235)

sqrt (chg x^2+chg y^2)=sqrt(49+529)=sqrt(578)=24.0